TD Auto Finance has 45,947 shares of stock outstanding at a market price of $18.8 a share. What will be the market price per share if the company does a 1-for-5 reverse stock split?
Answer #1

In a reverse stock split, the stocks are consolidated together

5 existing shares are merged into 1 new share

Hence, Number of new shares = 45947/5 = 9189

Market price of each share = 18.8*5 = $94